When a piece of metal mass of 28.4 g is dropped into a graduated cylinder containing 23.93 mL of water, the water level rises to 30.72 mL. What is the density of the metal?

Answer in units of g/cm^3


Do not round at all. Leave answer as is


1
Expert's answer
2021-10-13T04:54:21-0400

Volume of metal "=30.72-23.93=6.79ml"


Density "=\frac{28.4}{6.79}=4.18262150221g/cm^3"